# Uncertainty-Aware Surrogate-based Amortized Bayesian Inference for Computationally Expensive Models

### v1

@@ Appendix B Proofs
In UA-SABI, as per Eq. 8, we sample from the joint distribution
$p(\boldsymbol{\omega},\mathbf{y},\mathbf{c},\sigma\mid D_{T})=p(\mathbf{y}\mid\boldsymbol{\omega},\mathbf{c},\sigma)\,p(\boldsymbol{\omega})\,p(\mathbf{c},\sigma\mid D_{T})$ (14)
to train the conditional neural density estimator $q(\boldsymbol{\omega}\mid\mathbf{y})$. We only condition on the data $\mathbf{y}$ and treat the surrogate parameters $(\mathbf{c},\sigma)$ as nuisance parameters ignored by $q(\boldsymbol{\omega}\mid\mathbf{y})$. This means that we (implicitly) integrate over the distribution of $(\mathbf{c},\sigma)$. Hence, UA-SABI targets the following posterior:
[…]
On the other hand, E-Post targets the following posterior, as per Eq. 10:
[…]
which shows $p_{\rm UA-SABI}(\boldsymbol{\omega}\mid\mathbf{y},D_{T})=p_{\rm E-Post}(\boldsymbol{\omega}\mid\mathbf{y},D_{T})$. ∎
@@ C.1 Surrogate Models
In case study 1, we train a Bayesian PCE with 2-dimensional Legendre polynomials and a maximum total degree of 3, resulting in $D=10$ polynomials. We set a normal prior for the surrogate coefficients, $p(\mathbf{c})=\mathcal{N}(0,5)$, and a half-normal prior for the approximation error parameter, $p(\sigma)=\text{Half-}\mathcal{N}(0.5)$.
In case study 2, we consider a Bayesian PCE with 3-dimensional aPC polynomials (see Section 4.3) and a maximum total degree of $3$, resulting in $D=19$ polynomials. Following Bürkner et al. [2], we place a sparsity-inducing R2D2 prior on the surrogate coefficients $\mathbf{c}$ with $R^{2}\sim\text{Beta}(0.5,2)$. For the approximation error we set the prior as $p(\sigma)=\text{Half-}\mathcal{N}(0.1)$.
@@ C.2 Neural Posterior Estimation
For both case studies, we used the same NPE setup. The summary network $S_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{y})$ is a DeepSet [49] composed of two hidden layers, each containing 10 neurons. It outputs a 10-dimensional summary vector. For the inference network $I_{\boldsymbol{\varphi}}(\mathbf{s})$ we employ a coupling flow as implemented in Radev et al. [33]. The training process employed a cosine learning rate scheduler with an initial learning rate of $5\times 10^{-4}$ and a minimum learning rate fraction of $\alpha=10^{-6}$. The scheduler operated over a total of 12,800 steps, corresponding to 128 batches per epoch over 100 epochs. All NPE models were trained using the Adam optimizer [17] for 100 epochs.
@@ Appendix D Additional Results for Case Study 2: $\text{CO}_{2}$ Storage Model
In Fig. 7 we show that SABI and Point as well as UA-SABI and E-Post yield similar results in both revovery and calibration. The notably wide intervals observed for Point, especially compared to SABI in the recovery plots suggest convergence problems in the MCMC, likely due to the low standard deviation in the likelihood. These convergence problems may explain the inconsistencies between SABI and Point. In comparison, their uncertainty-aware counterparts UA-SABI and E-Post produce highly similar estimates.
Figure 7: Recovery plots comparing to ground truth and MCMC full reference solution, ECDF difference plots (from left to right) for $\text{CO}_{2}$ storage model porosity parameter for 200 ground truth samples: SABI vs. Point (top) and UA-SABI vs. E-Post (bottom). For ECDF difference plots, empirical ranks are shown in blue, $95\%$ confidence bands assuming calibration are shown in grey.

### v2

@@ Appendix B Proofs
In UA-SABI, as per Eq.˜8, we sample from the joint distribution
$p(\bm{\omega},\mathbf{y},\mathbf{c},\sigma\mid D_{T})=p(\mathbf{y}\mid\bm{\omega},\mathbf{c},\sigma)\,p(\bm{\omega})\,p(\mathbf{c},\sigma\mid D_{T})$ (14)
to train the conditional neural density estimator $q(\bm{\omega}\mid\mathbf{y})$. We only condition on the data $\mathbf{y}$ and treat the surrogate parameters $(\mathbf{c},\sigma)$ as nuisance parameters ignored by $q(\bm{\omega}\mid\mathbf{y})$. This means that we (implicitly) integrate over the distribution of $(\mathbf{c},\sigma)$. Hence, UA-SABI targets the following posterior:
$p_{\rm UA-SABI}(\bm{\omega}\mid\mathbf{y},D_{T})\propto\iint p(\mathbf{y}\mid\bm{\omega},\mathbf{c},\sigma)\,p(\bm{\omega})\,p(\mathbf{c},\sigma\mid D_{T})\,d\mathbf{c}\,d\sigma$ (15)
On the other hand, E-Post targets the following posterior, as per Eq.˜10:
$p_{\rm E-Post}(\bm{\omega}\mid\mathbf{y},D_{T})$ $=\iint p(\bm{\omega}\mid\mathbf{y},\mathbf{c},\sigma)\,p(\mathbf{c},\sigma\mid D_{T})\,d\mathbf{c}\,d\sigma$ (16)
$\propto\iint p(\mathbf{y}\mid\bm{\omega},\mathbf{c},\sigma)\,p(\bm{\omega})\,p(\mathbf{c},\sigma\mid D_{T})\,d\mathbf{c}\,d\sigma,$ (17)
which shows $p_{\rm UA-SABI}(\bm{\omega}\mid\mathbf{y},D_{T})=p_{\rm E-Post}(\bm{\omega}\mid\mathbf{y},D_{T})$. ∎
@@ C.1 Surrogate Models
In case study 1, we train a Bayesian PCE with 2-dimensional Legendre polynomials and a maximum total degree of 3, resulting in $J=10$ polynomials. We set a normal prior for the surrogate coefficients, $p(\mathbf{c})=\mathcal{N}(0,5)$, and a half-normal prior for the approximation error parameter, $p(\sigma)=\text{Half-}\mathcal{N}(0.5)$.
In case study 2 and 3, we consider a Bayesian PCE with 3- and 4-dimensional aPC polynomials for the $\text{CO}_{2}$ model (see Section˜4.3) and a maximum total degree of $3$, resulting in $J=19$ and $J=35$ polynomials respectively. Following Bürkner et al. (2023), we place a sparsity-inducing R2D2 prior on the surrogate coefficients $\mathbf{c}$ with $R^{2}\sim\text{Beta}(0.5,2)$. For the approximation error we set the prior as $p(\sigma)=\text{Half-}\mathcal{N}(0.1)$.
@@ C.2 Neural Posterior Estimation
For the first two case studies, we used the same NPE setup. The summary network $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ is a DeepSet Zaheer et al. (2017) composed of two hidden layers, each containing 10 neurons. It outputs a 10-dimensional summary vector. For the inference network $I_{\bm{\varphi}}(\mathbf{s})$, we employ a coupling flow as implemented in Radev et al. (2023b). For the third case study, we did not use a summary network as the measurement locations in the experiment were fixed. For the inference network $I_{\bm{\varphi}}(\mathbf{s})$, we used a spline coupling flow (Durkan et al., 2019) as implemented in Radev et al. (2023b). The training process employed a cosine learning rate scheduler with an initial learning rate of $5\times 10^{-4}$ and a minimum learning rate fraction of $\alpha=10^{-6}$. The scheduler operated over a total of 12,800 steps, corresponding to 128 batches per epoch over 100 epochs. All NPE models were trained using the Adam optimizer (Kingma & Ba, 2017) for 100 epochs.
@@ Appendix D Additional Results for Case Study 2: $\text{CO}_{2}$ Storage Model
In Fig.˜11 we show that SABI and Point as well as UA-SABI and E-Post yield similar results in both recovery and calibration. The notably wide intervals observed for Point, especially compared to SABI in the recovery plots suggest convergence problems in the MCMC, likely due to the low standard deviation in the likelihood. These convergence problems may explain the inconsistencies between SABI and Point. In comparison, their uncertainty-aware counterparts UA-SABI and E-Post produce highly similar estimates.
Figure 11: $\text{CO}_{2}$ recovery plots comparing to ground truth and MCMC full reference solution, ECDF difference plots (from left to right) for 200 ground truth samples: SABI vs. Point (top) and UA-SABI vs. E-Post (bottom). For ECDF difference plots, empirical ranks are shown in blue, $95\%$ confidence bands assuming calibration are shown in grey.
Appendix E Additional Results for Case Study 3: MICP Model
E.1 Parameter Posterior with Original Priors
Performing inference with the original prior for $k_{ub}$ leads, regardless of the inference method, to point-like posteriors at the upper boundary of the prior. The reason is the same for all methods: the maximal value of $k_{ub}$ yields the highest output values. Since the measurement data are always higher than the surrogate output under the original prior, the posterior collapses to a point-like estimate at the maximum admissible value of $k_{ub}$, as exemplified in Fig.˜8.
Figure 12: KDEs of 4,000 parameter posterior samples for $k_{ub}$ of the MICP model obtained with low-budget ABI, Point, E-Post, SABI and UA-SABI using the original parameter prior.
E.2 Sensitivity Analysis
We conducted a sensitivity analysis on the surrogate to identify the parameters to which the surrogate output is most sensitive. Sobol indices (Sobol’, 1990) decompose the output variance into contributions from each parameter (including their interactions), with the total Sobol index of a parameter quantifying its overall influence on the output variance. Figure˜13 presents the distributions of the total Sobol indices (Le Gratiet et al., 2017) for each parameter at each location, noting that each location has its own surrogate. Since the Bayesian PCE yields samples of the coefficients, the corresponding Sobol indices are also obtained in the form of distributions. Figure˜13 clearly shows that almost all the variance in the output at all locations is determined by $k_{ub}$.
Figure 13: Distribution of total Sobol indices over all locations for the four parameters $c_{a,1},c_{a,2},\rho_{f},k_{ub}$ of the MICP model.
E.3 Parameter Posterior for Remaining Parameters
Figure˜14 shows the KDEs of 4,000 parameter samples for the 3 remaining parameters, $c_{a,1},c_{a,2}$, and $\rho_{f}$, obtained with low-budget ABI, Point, E-Post, SABI, and UA-SABI. Two distinct scenarios emerge: 1) the posterior collapses to a point estimate (often at the boundary of the prior), or 2) the posterior remains close to the prior distribution. Point-like posteriors are produced by low-budget ABI or Point, while surrogate-based methods yield posteriors resembling the prior. Both outcomes can be explained by the lack of sensitivity of the output to these parameters, as shown by the total Sobol indices in Fig.˜13. This missing sensitivity makes reliable inference very difficult.
Figure 14: KDEs of 4,000 parameter posterior samples for the remaining parameters $c_{a,1},c_{a,2}$, and $\rho_{f}$ of the MICP model obtained with low-budget ABI, Point, E-Post, SABI and UA-SABI.
E.4 Runtime Comparison
To compare runtimes as in case studies 1 and 2, inference must be performed on multiple datasets. Since only one dataset of real measurements is available, we conducted the runtime analysis on synthetic datasets. The experiments were also run on a computing cluster with two AMD EPYC 7551 CPUs (totaling 64 physical cores) to speed up E-Post through parallelization.
Figure˜15 shows the measured runtimes of UA-SABI and E-Post for 2,3,4,5{2,3,4,5} inference runs, with E-Post parallelized across 16 cores. We observe a break-even point between 3 and 4 runs, meaning that UA-SABI becomes advantageous starting from 4 runs. Despite E-Post leveraging more cores, the training costs of UA-SABI are amortized earlier when the model is more expensive.
Figure 15: Comparison of runtimes for MICP for {2,3,4,5} runs for UA-SABI (training and inference) and E-Post (inference) with break-even point.